Solve the following equations and check your results.2x – 1 = 14 – x 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2x–1=14–x
⇒2x+x=14+1
⇒3x=15
⇒x=5
Putting the value of x in RHS and LHS we get,
(2×5) - 1 = 14 - 5
⇒10-1=9
⇒9=9
⇒LHS=RHS

Thanks